Determine the energy of a photon whose frequency is 5.87 × 1017 Hz.
Answer in units of J.
1
Expert's answer
2020-02-21T07:53:27-0500

Energy of a photon is calculated by the formula:

"E=h\nu" , where "h=6.63*10^{-34}" J*s;

"E=6.63*10^{-34}*5.87*10^{17}=3.892*10^{-16}" J.


Answer: 3.892*10-16 J.